### v2

# Stochastic Subspace via Probabilistic Principal Component Analysis for Characterizing Model Error

Abstract
This paper proposes a probabilistic model of subspaces based on the probabilistic principal component analysis (PCA). Given a sample of vectors in the embedding space—commonly known as a snapshot matrix—this method uses quantities derived from the probabilistic PCA to construct distributions of the sample matrix, as well as the principal subspaces. It is applicable to projection-based reduced-order modeling methods, such as proper orthogonal decomposition and related model reduction methods. The stochastic subspace thus constructed can be used, for example, to characterize model-form uncertainty in computational mechanics. The proposed method has multiple desirable properties: (1) it is naturally justified by the probabilistic PCA and has analytic forms for the induced random matrix models; (2) it satisfies linear constraints, such as boundary conditions of all kinds, by default; (3) it has only one hyperparameter, which significantly simplifies training; and (4) its algorithm is very easy to implement. We demonstrate the performance of the proposed method via several numerical examples in computational mechanics and structural dynamics.
Keywords Model error  $\cdot$ Model-form uncertainty  $\cdot$ Stochastic reduced-order modeling  $\cdot$ Probabilistic principal component analysis  $\cdot$ Stochastic proper orthogonal decomposition
1 Introduction
Engineering systems can often be described by a set of partial differential equations, commonly referred to as governing equations. These governing equations are typically derived by making assumptions and simplifying the underlying physical process, which inevitably introduces errors. The errors are further manifested via numerical approximations, imprecise initial and boundary conditions, unknown model parameters, and errors in the differential equations themselves. Model error (also known as model discrepancy, model inadequacy, structural uncertainty, or model-form error in various contexts) is ubiquitous in computational engineering, but its probabilistic analysis has been a notoriously difficult challenge. Unlike parametric uncertainties that are associated with model parameters, model-form uncertainties concern the variability of the model itself, which are inherently nonparametric. Pernot [1] argues that model-form uncertainty should not be absorbed in parametric uncertainty as it leads to prediction bias. The analysis of model-form uncertainty can be divided into two sub-problems: 1) characterization and 2) correction. Determining the predictive error of an inadequate model is categorized as characterization. Reducing the predictive error of an inadequate model via some form of adjustment is called correction, which is much more complex than characterization. The central focus of our current work is to characterize model error. There are various studies in the literature that address model error. The approaches can be divided into two categories: 1) direct representation and 2) indirect representation.
The methods based on direct representation are mainly subdivided into two categories, depending on where model errors are addressed: 1) external, and 2) internal [2]. In external direct representation methods, model error is accounted for by adding a correction term to the model output, which is usually calibrated using a Gaussian process to match the observed data. One of the earliest studies to address model error was carried out by Kennedy and O’Hagan [3] in 2001, in which the authors presented a Bayesian calibration technique as an improvement on the traditional techniques. Since then, the Bayesian inferential and modeling framework has been adopted and further developed by many studies [4, 5, 6, 7, 8]. KOH Bayesian calibration corrects the model output but has limited extrapolative capability to unobserved quantities [9]. Farrell et al. [10] present an adaptive modeling paradigm for Bayesian calibration, and model selection, where the authors developed an Occam-Plausibility algorithm to select the best model with calibration. However, if the initial set of models is invalid, then another set of models is required, which requires a lot of prior information and modeling experience. In general, methods using external direct representation tend to violate physical constraints [11, 12], rely heavily on prior information [13], fail to elucidate model error and data error, and struggle with extrapolation and prediction of unobserved quantity [6, 14, 1]. To alleviate some of the drawbacks, Brynjarsdóttir and O’Hagan [13] and Plumlee [15] show that prior does play an important role in improving performance. He and Xiu [2] proposed a general physics-constrained model correction framework, ensuring that the corrected model adheres to physical constraints. Despite all the developments, methods based on external direct representation have several limitations, as listed above, mainly because model error is inherently nonparametric.
Internal direct representation methods improve a model through intrusive modifications, rather than by adding an extra term to the output. Embedded model enrichment [14] provides a framework for extrapolative predictions using composite models. Morrison et al. [16] address model inadequacy with stochastic operators introduced as a source in the governing equations while preserving underlying physical constraints. Similarly, Portone and Moser [17] propose a stochastic operator for an unclosed dispersion term in an averaged advection–diffusion equation for model closure. However, modifying the governing equations of the model brings challenges in the computational cost and implementation. Additionally, significant prior knowledge of the system is required to alter the governing equations and achieve improvements effectively. Bhat et al. [18] address model-form error by embedding dynamic discrepancies within small-scale models, using Bayesian smoothing splines (BS-ANOVA) framework. Strong and Oakley [11] introduce a novel way to quantify model inadequacies by decomposing the model into sub-functions and addressing discrepancies in each sub-function. The authors show the application of their method in health economics studies. Sargsyan et al. [19] improve uncertainty representation and quantification in computational models by introducing a Bayesian framework with embedded model error representation using polynomial chaos expansion. Other studies have used internal model error correction for various application areas, e.g., Reynolds-averaged Navier Stokes (RANS) applications [20], large eddy simulation (LES) computations [21], molecular simulations [22], particle-laden flows [23], and chemical ignition modeling [24]. In general, internal direct representation methods require a deep understanding of the system to implement the changes necessary for improvement. This knowledge is not always available and often demands an extensive study of the system.
An approach based on indirect representation is presented by Sargsyan et al. [12], where the authors aim to address model error at the source by hyperparameterizing the parameters of assumptions. However, it is not always easy to pinpoint the source of the errors especially when errors arise from multiple sources. Furthermore, this approach does not correct for model error. Another indirect representation approach to address model error is based on the concept of stochastic reduced order modeling, originally introduced by Soize in the early 2000s [25, 26, 27] and reviewed in [28]. In their earlier approaches, the authors build the reduced-order model (ROM) via projection onto a deterministic subspace and construct stochastic ROMs by randomizing the coefficient matrices using the principle of maximum entropy. The term “stochastic reduced order model” is also used by a method developed around 2010 in a different context [29, 30, 31], which is essentially a discrete approximation of a probability distribution and can be constructed non-intrusively from a computational model.
A recent work by Soize and Farhat [32] in model-form uncertainty merges ideas from projection-based reduced-order modeling [33, 34] and random matrix theory [35, 36]. In the new approach, instead of randomizing coefficient matrices, the reduced-order basis (ROB) is randomized. The key observation is that since the ROB determines the ROM, randomizing the ROB also randomizes the ROM, which can be used for efficient probabilistic analysis of model error. Using computationally inexpensive ROM instead of a full-order model makes the approach computationally tractable for uncertainty quantification tasks. The probabilistic model and estimation procedure proposed in [32] have been applied to various engineering problems [37]. Because this method can face challenges due to a large number of hyperparameters and complex implementation, later developments have aimed at reducing the number of hyperparameters and simplifying hyperparameter training [38]. In settings with multiple candidate models, Zhang and Guilleminot [39] proposed a different probabilistic model for ROBs, and has been used for various problems in computational mechanics [40, 41].
This paper proposes a novel framework for probabilistic modeling of principal subspaces. The stochastic ROMs constructed from such stochastic subspaces enable improved characterization of model uncertainty in computational mechanics. The main contributions of our work are as follows:
  1. 1.
    We introduce a new class of probabilistic models of subspaces, which is simple, has nice analytical properties and can be sampled efficiently. This is of general interest, and can be seen as a stochastic version of the proper orthogonal decomposition.
  2. 2.
    We use this stochastic subspace model to construct stochastic ROMs, and reduce the number of hyperparameters to a single scalar. This contrasts with existing stochastic ROM approaches, which use stochastic bases and typically involve more hyperparameters.
  3. 3.
    The optimization of the hyperparameter is fully automated and done efficiently by exploiting one-dimensional optimization algorithms.
  4. 4.
    We demonstrate in a series of numerical experiments that our method provides consistent and sharp uncertainty estimates, with low computational costs.
This paper is organized as follows: A brief introduction to stochastic reduced-order modeling is presented in Section 2. The proposed stochastic subspace model, along with the algorithm, is described in Section 3. Critical information regarding the hyperparameter optimization is given in Section 4. Related works in the literature are reviewed in Section 5. The accuracy and efficiency of the proposed method are validated using numerical examples in Section 6. Finally, Section 7 concludes the paper with a brief summary and potential future work.
2 Stochastic reduced-order modeling
This section presents the concept of stochastic reduced-order modeling. The high-dimensional model (also known as the full-order model) is described in Section 2.1. Section 2.2 describes the reduced-order model and its formulation from a high-dimensional model. Finally, Section 2.3 presents the stochastic reduced-order model and its formulation from a reduced-order model. The presentation in this section is very general and can easily be applied to any engineering system.
2.1 High-dimensional model
In general, we consider a parametric nonlinear system given by a set of ordinary differential equations (ODEs):
$\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x},t;\boldsymbol{\mu}),$ (1)
with $\mathbf{x}\in\mathbb{R}^{n}$ and $t\in[0,\infty)$, and is subject to initial conditions $\mathbf{x}(0)=\mathbf{x}_{\_}0$ and linear constraints (e.g., boundary conditions) $\mathbf{B}^{\intercal}\mathbf{x}=0$, where $\mathbf{B}\in\mathbb{R}^{n\times n_{\_}{CD}}$. These equations often come from a spatial discretization of a set of partial differential equations that govern a given physical system, with dimension $n\gg 1$. We call this the high-dimensional model (HDM). If the system is time-independent, Eq. 1 reduces to a set of algebraic equations: $\mathbf{f}(\mathbf{x};\boldsymbol{\mu})=0$.
2.2 Reduced-order model
The Galerkin projection of the HDM onto a $k$-dim subspace $\mathscr{V}$ of the state space $\mathbb{R}^{n}$ gives a reduced-order model (ROM), which can be solved faster than the HDM. Let $\mathbf{V}\in\text{St}(n,k)$ be an orthonormal basis of the subspace $\mathscr{V}$, then the ROM can be written as:
$\mathbf{x}=\mathbf{V}\mathbf{q},\quad\dot{\mathbf{q}}=\mathbf{V}^{\intercal}\,\mathbf{f}(\mathbf{V}\mathbf{q},t;\boldsymbol{\mu}),$ (2)
with $\mathbf{q}\in\mathbb{R}^{k}$ and initial conditions $\mathbf{q}(0)=\mathbf{V}^{\intercal}\mathbf{x}_{\_}0$. To satisfy the linear constraints, we must have $\mathbf{B}^{\intercal}\mathbf{V}=0$. For time-independent systems, Eq. 2 reduces to a set of algebraic equations: $\mathbf{V}^{\intercal}\,\mathbf{f}(\mathbf{V}\mathbf{q};\boldsymbol{\mu})$ = 00.
2.3 Stochastic reduced-order model
The stochastic reduced-order model (SROM) builds upon the structure of ROM with the difference that the deterministic ROB $\mathbf{V}$ is replaced by its stochastic counterpart $\mathbf{W}$. The change from a deterministic basis to a stochastic basis introduces randomness to the model and allows it to capture model error. The SROM can be written as:
$\mathbf{x}=\mathbf{W}\mathbf{q},\quad\dot{\mathbf{q}}=\mathbf{W}^{\intercal}\,\mathbf{f}(\mathbf{W}\mathbf{q},t;\boldsymbol{\mu}),\quad\mathbf{W}\sim\mu_{\_}{\mathbf{V}}$ (3)
where stochastic basis $\mathbf{W}$ follows a probability distribution $\mu_{\_}{\mathbf{V}}$. However, it is more natural to impose a probability distribution $\mu_{\_}{\mathscr{V}}$ on the subspace rather than the basis, because the Galerkin projection is uniquely determined by the subspace and is invariant under changes of basis. In this work, we use the stochastic subspace model introduced in Section 3 to sample stochastic basis $\mathbf{W}$ and construct SROM, which is then used to characterize model error.
3 Stochastic subspace model
This section presents the formulation of the stochastic subspace model using probabilistic principal component analysis. The derivation of a deterministic principal subspace via proper orthogonal decomposition is outlined in Section 3.1. The procedure of constructing random matrices is described in Section 3.2. The stochastic subspace model and the sampling algorithm are discussed in Section 3.3.
3.1 Subspace from the PCA: proper orthogonal decomposition
Principal component analysis (PCA) is a common technique for dimension reduction. Closely related to PCA, a common way to find the ROB $\mathbf{V}$ and the associated subspace $\mathscr{V}$ is the proper orthogonal decomposition (POD). Following the notations in Section 2, let $\mathbf{X}=[\mathbf{x}_{\_}1\,\cdots\,\mathbf{x}_{\_}m]\in\mathbb{R}^{n\times m}$ be a sample of the state, with $\mathbf{x}_{\_}i=\mathbf{x}(t_{\_}i;\boldsymbol{\mu}_{\_}i)$, sample mean $\overline{\mathbf{x}}=\frac{1}{m}\sum_{\_}{i=1}^{m}\mathbf{x}_{\_}i$, centered sample $\mathbf{X}_{\_}0=\mathbf{X}-\overline{\mathbf{x}}\mathbf{1}_{\_}m^{\intercal}$, and sample covariance $\mathbf{S}=\frac{1}{m}\mathbf{X}_{\_}0\mathbf{X}_{\_}0^{\intercal}$. Let $\mathbf{X}_{\_}0=\mathbf{V}_{\_}r\operatorname{diag}(\boldsymbol{\sigma}_{\_}r)\mathbf{W}_{\_}r^{\intercal}$ be a compact singular value decomposition (SVD), where $\boldsymbol{\sigma}_{\_}r\in\mathbb{R}^{r}_{\_}{>0\downarrow}$ is in non-increasing order. The sample covariance matrix $\mathbf{S}$ can be written as $\mathbf{S}=\mathbf{V}_{\_}r\operatorname{diag}(\boldsymbol{\sigma}_{\_}r/\sqrt{m})^{2}\mathbf{V}_{\_}r^{\intercal}$. POD takes the principal basis $\mathbf{V}_{\_}k$—the leading $k$ eigenvectors of $\mathbf{S}$—as the deterministic ROB. The corresponding subspace is the principal subspace $\mathscr{V}_{\_}k:=\text{range}(\mathbf{V}_{\_}k)$. Since all state samples satisfy the linear constraints, $\mathbf{V}_{\_}k$ satisfies the constraints automatically. POD has several other desirable properties: it can extract coherent structures, provide error estimates, and is computationally efficient and straightforward to implement [42].
3.2 Random matrices from the PPCA
Probabilistic PCA (PPCA) [43] reformulates PCA as the maximum likelihood solution of a probabilistic latent variable model. Here, we briefly overview PPCA and introduce some related random matrix models.
Assume that the data-generating process can be modeled as: $\mathbf{x}=\boldsymbol{\mu}+\mathbf{U}\mathbf{z}+\boldsymbol{\epsilon}$, where $\boldsymbol{\mu}\in\mathbb{R}^{n}$, $\mathbf{U}\in\mathbb{R}^{n\times k}$, $\mathbf{z}\sim N_{\_}k(0,\mathbf{I}_{\_}k)$, $\boldsymbol{\epsilon}\sim N_{\_}n(0,\varepsilon^{2}\mathbf{I}_{\_}n)$, $n\in\mathbb{Z}_{\_}{>0}$ is the ambient dimension, and $k\in\{1,\cdots,n\}$ is the latent dimension. Then the observed data follows a Gaussian distribution $\mathbf{x}\sim N_{\_}n(\boldsymbol{\mu},\mathbf{C})$, where covariance matrix $\mathbf{C}=\mathbf{U}\mathbf{U}^{\intercal}+\varepsilon^{2}\mathbf{I}_{\_}n$.
Assume that we have a data sample $\mathbf{X}=[\mathbf{x}_{\_}1\,\cdots\,\mathbf{x}_{\_}m]\in\mathbb{R}^{n\times m}$ of size $m$. Let $\mathbf{S}=\mathbf{V}\operatorname{diag}(\boldsymbol{\lambda})\mathbf{V}^{\intercal}$ be an eigenvalue decomposition (EVD) of the sample covariance matrix, where $\mathbf{V}\in O(n)$ is an orthogonal matrix and $\boldsymbol{\lambda}\in\mathbb{R}^{n}_{\_}{\geq 0\downarrow}$ is in non-increasing order. The maximum likelihood estimator of the model parameters $(\boldsymbol{\mu},\mathbf{U},\varepsilon^{2})$ are: $\widetilde{\boldsymbol{\mu}}=\overline{\mathbf{x}}$, $\widetilde{\mathbf{U}}=\mathbf{V}_{\_}k[\operatorname{diag}(\lambda_{\_}i-\tilde{\varepsilon}^{2})_{\_}{i=1}^{k}]^{1/2}\mathbf{Q}$, and $\tilde{\varepsilon}^{2}=\frac{1}{n-k}\sum_{\_}{i=k+1}^{n}\lambda_{\_}i$, where $\mathbf{V}_{\_}k=[\mathbf{v}_{\_}1\,\cdots\,\mathbf{v}_{\_}k]$ consists of the first $k$ columns of $\mathbf{V}$, and $\mathbf{Q}\in O(k)$ is any order-$k$ orthogonal matrix. The estimated covariance matrix can thus be written as $\widetilde{\mathbf{C}}=\mathbf{V}_{\_}k\operatorname{diag}(\lambda_{\_}i-\tilde{\varepsilon}^{2})_{\_}{i=1}^{k}\mathbf{V}_{\_}k^{\intercal}+\tilde{\varepsilon}^{2}\mathbf{I}_{\_}n$. We call the columns of $\mathbf{V}_{\_}k$ principal components of the data sample and $\boldsymbol{\lambda}_{\_}k=(\lambda_{\_}i)_{\_}{i=1}^{k}$ the corresponding principal variances.
The PPCA assumes that the latent dimension $k$ is known, and estimates the unknown data distribution $\mathbf{x}\sim N_{\_}n(\boldsymbol{\mu},\mathbf{C})$ by maximum likelihood as $\widetilde{\mathbf{x}}\sim N_{\_}n(\widetilde{\boldsymbol{\mu}},\widetilde{\mathbf{C}})$. Without assuming $k$ as known, the data distribution can simply be modeled as $\widetilde{\mathbf{x}}\sim N_{\_}n(\overline{\mathbf{x}},\mathbf{S})$, the multivariate Gaussian distribution parameterized by the sample mean and the sample covariance. From this empirical data distribution, we can derive distributions of some related matrices, all in analytical form.
[…]
$\widetilde{\mathbf{V}}_{\_}k:=\pi(\widetilde{\mathbf{X}}_{\_}{0,k})\sim\text{MACG}_{\_}{n,k}(\mathbf{S}),$ (4)
where $\widetilde{\mathbf{X}}_{\_}{0,k}\sim N_{\_}{n,k}(0;\mathbf{S},\mathbf{I}_{\_}k)$ is a centered $k$-sample matrix and $\pi(\mathbf{M}):=\mathbf{M}(\mathbf{M}^{\intercal}\mathbf{M})^{-1/2}$ is orthonormalization by polar decomposition. As a mapping, $\pi:\mathbb{R}^{n\times k}_{\_}*\mapsto\text{St}(n,k)$ is uniquely defined for all full column-rank $n$-by-$k$ matrices, $k\leq n$, and takes values in the Stiefel manifold $\text{St}(n,k):=\{\mathbf{V}\in\mathbb{R}^{n\times k}:\mathbf{V}^{\intercal}\mathbf{V}=\mathbf{I}_{\_}k\}$. The range (i.e., column space) of the data sample matrix of size $k$ also has an MACG distribution:
$\widetilde{\mathscr{V}}_{\_}k:=\text{range}(\widetilde{\mathbf{V}}_{\_}k)\sim\text{MACG}_{\_}{n,k}(\mathbf{S}),$ (5)
which is supported on the Grassmann manifold $\text{Gr}(n,k)$ of $k$-dim subspaces of the Euclidean space $\mathbb{R}^{n}$. Its probability density function (PDF) has a unique mode—and therefore a unique global maximal point—at the principal subspace $\mathscr{V}_{\_}k$, see A. This distribution is useful, for example, for modeling subspaces for parametric ROM [34].
3.3 Stochastic subspace via the PPCA
Here we propose a new class of stochastic subspace models $\text{MACG}_{\_}{n,k,\beta}(\mathbf{S})$ with $\beta\in[k,\infty)$. The classical $\text{MACG}_{\_}{n,k}(\mathbf{S})$ distribution is a special case with $\beta=k$.
Define the principal subspace map $\pi_{\_}k(\mathbf{X}):=\text{range}(\mathbf{U}_{\_}k)$, where $\mathbf{U}_{\_}k$ consists of singular vectors associated with the $k$ largest singular values of $\mathbf{X}$. As a mapping, $\pi_{\_}k:\mathbb{R}^{n\times m}_{\_}{k>}\mapsto\text{Gr}(n,k)$ is uniquely defined for all $n$-by-$m$ matrices whose $k$-th largest singular value is larger than the $(k+1)$-th, with $k\leq\min(n,m)$. Using this map, the POD subspace can be written as: $\mathscr{V}_{\_}k=\pi_{\_}k(\mathbf{X}_{\_}0)$. We define a stochastic subspace model:
[…]
where $\widetilde{\mathbf{X}}_{\_}{0,\beta}\sim N_{\_}{n,\beta}(0;\mathbf{S},\mathbf{I}_{\_}{\beta})$ is a centered $\beta$-sample matrix and $\beta\in\{k,k+1,\cdots\}$ is the resample size. Considering its close connection with the POD, we may call this model a stochastic POD. Similar to $\text{MACG}_{\_}{n,k}$, the unique mode and global maximal of $\text{MACG}_{\_}{n,k,\beta}$ is the principal subspace $\mathscr{V}_{\_}k$. The hyperparameter $\beta$ controls the concentration of the distribution: larger value means less variation around $\mathscr{V}_{\_}k$. As with POD, the stochastic basis $\mathbf{W}$ sampled from the stochastic subspace model defined in Eq. 6 satisfies the linear constraints $\mathbf{B}^{\intercal}\mathbf{W}=0$ automatically.
Sampling from this model can be done very efficiently for $n\gg 1$ if $k$ and the rank of $\mathbf{S}$ are small, which is often the case in practical applications. Let $\mathbf{S}=\mathbf{V}_{\_}r\operatorname{diag}(\boldsymbol{\lambda}_{\_}r)\mathbf{V}_{\_}r^{\intercal}$ be a compact EVD where $r=\text{rank}(\mathbf{S})$. We can show that:
$\mathscr{W}=\mathbf{V}_{\_}r\,\pi_{\_}k\left(\operatorname{diag}(\boldsymbol{\lambda}_{\_}r)^{1/2}\,\mathbf{Z}_{\_}{r\times\beta}\right),$ (7)
where $\mathbf{Z}_{\_}{r\times\beta}$ is an $r$-by-$\beta$ standard Gaussian matrix. Algorithm 1 gives a procedure for sampling $\text{MACG}_{\_}{r,k,\beta}(\mathbf{S})$ with $\mathbf{S}=\operatorname{diag}(\mathbf{s})^{2}$, which in effect implements the map $\pi_{\_}k(\operatorname{diag}(\mathbf{s})\,\mathbf{Z}_{\_}{r\times\beta})$. For a general covariance matrix $\mathbf{S}$, an orthonormal basis of a random subspace sampled from $\text{MACG}_{\_}{n,k,\beta}(\mathbf{S})$ can be obtained by left multiplying the output of Algorithm 1 with $\mathbf{V}_{\_}r$. Using Eq. 7 instead of Eq. 6 reduces the computational cost for truncated SVD from $O(nk\beta)$ to $O(rk\beta)$.
Algorithm 1 SS-PPCA: Stochastic subspace via probabilistic principal component analysis.
1:scale vector $\mathbf{s}\in\mathbb{R}^{r}_{\_}{>0}$; subspace dimension $k\in\{1,\cdots,r\}$; resample size $\beta\in\{k,k+1,\cdots\}$.
[…]
3:$\mathbf{M}\leftarrow\operatorname{diag}(\mathbf{s})\,\mathbf{Z}_{\_}{r\times\beta}$
4:Truncated SVD: $[\mathbf{U}_{\_}k,\mathbf{d}_{\_}k,\mathbf{V}_{\_}k]\leftarrow\text{svd}(\mathbf{M},k)$
5:$\mathbf{W}=\mathbf{U}_{\_}k$, an orthonormal basis of a random subspace sampled from $\text{MACG}_{\_}{r,k,\beta}(\mathbf{S})$ with $\mathbf{S}=\operatorname{diag}(\mathbf{s})^{2}$.
We can generalize the concentration hyperparameter $\beta$ to real numbers in $[k,\infty)$, which leads to a continuous family of distributions $\text{MACG}_{\_}{n,k,\beta}(\mathbf{S})$ and can be useful when the optimal value of $\beta$ is not much greater than one. In this case, we sample $\mathbf{Z}\in\mathbb{R}^{r\times\lceil\beta\rceil}$, where $\lceil\beta\rceil$ denotes the smallest integer greater than or equal to $\beta$. To account for the effect of real-valued $\beta$, the weight of the final column of $\mathbf{Z}$ is set to $\beta-\lfloor\beta\rfloor$, where $\lfloor\beta\rfloor$ is the largest integer less than or equal to $\beta$. The rest of the steps are the same as Algorithm 1.
4 Hyperparameter training
To find the optimal hyperparameter $\beta\in[k,\infty)$, we minimize the following objective function:
$f(\beta):=\mathbb{E}[|d_{\_}o(\mathbf{u}_{\_}L)-d_{\_}o(\mathbf{u}_{\_}E)|^{2}\,|\,\beta],$ (8)
where $\mathbf{u}_{\_}E$ is the experimental or ground-truth observation of the output, $\mathbf{u}_{\_}L$ is the low-fidelity prediction of the SROM, and $d_{\_}o(\mathbf{u}):=\|\mathbf{u}-\mathbf{u}_{\_}L^{o}\|_{\_}{L^{2}}$ is the $L^{2}$ distance to the low-fidelity prediction $\mathbf{u}_{\_}L^{o}$ of a reference model. Given the value of the hyperparameter $\beta$, the stochastic subspace model determines the stochastic ROM, which produces stochastic predictions $\mathbf{u}_{\_}L$ that is then summarized into the random variable $d_{\_}o(\mathbf{u}_{\_}L)$. Given the experimental or ground-truth observation $\mathbf{u}_{\_}E$, the objective function measures the mean squared error between $d_{\_}o(\mathbf{u}_{\_}L)$ and $d_{\_}o(\mathbf{u}_{\_}E)$, which is a statistical measure of how closely the SROM resembles the ground truth. Overall, this optimization problem aims to improve the consistency of the SROM in characterizing the error of the reference model.
We optimize the objective function $f(\beta)$ efficiently using a one-dimensional optimization algorithm, such as golden section search and successive parabolic interpolation. Implementations of such algorithms are readily available in many programming languages, e.g., fminbnd() in Matlab, scipy.optimize.minimizescalar() in Python, and optimize() in R. The objective function, however, is not accessible due to the expectation over SROM. To reduce computation, we approximate $f(\beta)$ at integer $\beta$ values by Monte Carlo sampling, using 1,000 random samples of the SROM, and linearly interpolate $f(\beta)$ between consecutive integer points. In practice, when the optimization algorithm queries $f(\beta)$ at a real-valued $\beta$, its values at the two closest integer points will be computed using Monte Carlo sampling and stored. If future queries need the value of $f(\beta)$ at a previously evaluated integer point, the stored value will be used to avoid re-evaluation.
The optimization scheme employed in this study is very easy to implement, although it may not be the most efficient due to Monte Carlo approximation of the expectation operator. We leave improvements to the hyperparameter optimization procedure for future work.
5 Related works
This section reviews in detail the two main works that use SROM for model error characterization. Firstly, the work by Soize and Farhat [32] is discussed in Section 5.1. In Section 5.2, the work by Zhang and Guilleminot [39] for characterizing model-form uncertainty across multiple models is presented. In Section 5.3, we distinguish the three methods based on their operational mechanisms and their dependence on available information.
5.1 Non-parametric model (NPM)
The seminal work of NPM [32] introduced the idea of using SROM to analyze model-form uncertainty. The stochastic basis in the NPM can be written as:
$\mathbf{W}=\pi\left(\overline{\mathbf{V}}+\mathcal{P}^{\text{St}}_{\_}{\overline{\mathbf{V}}}(\mathbf{P}_{\_}{\mathscr{B}^{\perp}}\mathbf{U})\right),\quad\mathbf{U}=s\mathbf{G}\mathbf{R},\quad\mathbf{G}\sim\mathcal{GP}(\gamma).$ (9)
The random matrix $\mathbf{G}$ requires knowledge of the discretization $\mathscr{D}$ of the spatial fields, and is sampled from a Gaussian process with a hyperparameter $\gamma$ for correlation length-scale. The column dependence matrix $\mathbf{R}$ is an order-$k$ upper triangular matrix with positive diagonal entries. The scale $s\in[0,\infty)$ is a non-negative number. Once $\mathbf{U}$ is constructed, orthogonal projection $\mathbf{P}_{\_}{\mathscr{B}^{\perp}}$ enforces the linear constraints, tangential projection $\mathcal{P}^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ projects a matrix to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$, and orthonormalization $\pi$ as in Eq. 4 gives an orthonormal basis. The hyperparameters $(s,\gamma,\mathbf{R})$ have a dimension of $k(k+1)/2+2$, and are trained by minimizing a weighted average of $J_{\_}{\text{mean}}$ and $J_{\_}{\text{std}}$: the former aims at matching the SROM mean prediction to ground-truth observations $\mathbf{O}$; the latter aims at matching the SROM standard deviation to a scaled difference between $\mathbf{O}$ and $\mathbf{u}_{\_}L^{o}$.
5.2 Riemannian stochastic model (RSM)
For problems where there exist a number of physically plausible candidate models, the RSM [39] characterizes model-form uncertainty using SROMs constructed with the following stochastic basis:
$\mathbf{W}=\exp^{\text{St}}_{\_}{\overline{\mathbf{V}}}\Big{\{}c\sum_{\_}{i=1}^{q}p_{\_}i\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}(\mathbf{V}^{(i)})\Big{\}},\quad\mathbf{p}\sim\text{Dirichlet}(\boldsymbol{\alpha}),$ (10)
where $c\in[0,\infty)$ is a scale parameter, $q$ is the number of models, and $\mathbf{p}=(p_{\_}i)_{\_}{i=1}^{q}$ is a probability vector following the Dirichlet distribution with concentration parameters $\boldsymbol{\alpha}\in\mathbb{R}_{\_}{>0}^{q}$. The Riemannian logarithm $\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a model-specific ROB $\mathbf{V}^{(i)}$ to the tangent space of the Stiefel manifold at a reference ROB $\overline{\mathbf{V}}$. The Riemannian exponential $\exp^{\text{St}}_{\_}{\overline{\mathbf{V}}}$ maps a tangent vector at $\overline{\mathbf{V}}$ back to the Stiefel manifold, giving the orthonormal basis $\mathbf{W}$. The scale parameter $c$ is usually set to one or omitted in later works [40, 41]. The hyperparameters $\boldsymbol{\alpha}$ are trained by minimizing $\|\mathbb{E}[\log^{\text{St}}_{\_}{\overline{\mathbf{V}}}\mathbf{W}]\|_{\_}F^{2}$, to match the center of mass of the distribution to $\overline{\mathbf{V}}$. It is a quadratic program and can be solved efficiently. We note that the RSM cannot be applied outside multi-model settings.
[…]
Figure 1: Dependence diagrams for (from top to bottom) the NPM and the RSM models of stochastic basis and the SS-PPCA model of stochastic subspace. The dashed lines indicate that the connection only exists when characterizing the ROM-to-HDM error.
5.3 Comparison of model structures
To deepen understanding of the three methods, we outline their structures in Figure 1 and highlight a few key differences below. Both NPM and RSM construct probabilistic models of reduced-order bases, whereas SS-PPCA models subspaces directly. Another distinction lies in the reliance on discretization information: NPM requires discretization and boundary condition data for hyperparameter optimization and SROM sampling, while SS-PPCA and RSM do not. Finally, while RSM optimizes its hyperparameters using only the reduced bases, SS-PPCA and NPM incorporate observed quantities of interest into their objective functions.
6 Numerical experiments
In this section, we evaluate the proposed method, SS-PPCA, in characterizing model error using three numerical examples. In Section 6.1, we examine a parametric linear static problem and characterize model error at unseen parameters. In Section 6.2, we demonstrate how our approach can be used to characterize the HDM error, assuming that experimental data is available. Finally, in Section 6.3, we discuss a linear dynamics problem involving a space structure. All data and code for these examples are available at https://github.com/UQUH/SSPPCA.
The performance of the SROMs is assessed using two general metrics for stochastic models: consistency and sharpness. By consistent UQ, we mean that the statistics on prediction uncertainty—such as standard deviation and predictive intervals (PI)—derived from the model match the prediction error on average. Without consistency, the model either under- or over-estimates its prediction error, which risks being over-confident or unnecessarily conservative. By sharp UQ, we mean the model’s PIs are narrower than alternative methods. This is essentially the probabilistic version of model accuracy and is desirable as we want to minimize error.
6.1 Parametric linear static problem
We consider a one-dimensional parametric linear static problem with $n=$1,000$$ degrees of freedom (DoFs), governed by the system of equations:
$\mathbf{K}\mathbf{x}(\boldsymbol{\mu})=\mathbf{f}(\boldsymbol{\mu}),$ (11)
[…]
Figure 2: HDM vs ROM displacement at the test parameter
In this example, $\mathbf{K}$ does not depend on $\boldsymbol{\mu}$. Hence, it can be factorized once and stored for efficient inversion. However, for a general case, the parametric HDM can be computationally expensive, especially if it needs to be evaluated multiple times. This justifies the use of ROM. We construct a ROM by evaluating the HDM at 100 parameter samples, $\boldsymbol{\mu}_{\_}j\in[0,1]^{5}$, drawn via Latin Hypercube Sampling (LHS). For each sample, the solution is computed as $\mathbf{x}_{\_}j=\mathbf{x}(\boldsymbol{\mu}_{\_}j)=\mathbf{K}^{-1}\mathbf{f}(\boldsymbol{\mu}_{\_}j)$, $j=1,\cdots,100$. These solutions are concatenated to form a snapshot matrix $\mathbf{X}$, which is used to construct a deterministic ROB $\mathbf{V}=\pi_{\_}k(\mathbf{X})$ with dimension $k=4$ via POD. The governing equation of the ROM is given by:
$\mathbf{x}_{\_}{R}=\mathbf{V}\mathbf{q},\quad\mathbf{K}_{\_}k\mathbf{q}=\mathbf{V}^{\intercal}\mathbf{f}(\mu),\quad\mathbf{K}_{\_}k=\mathbf{V}^{\intercal}\mathbf{K}\mathbf{V}.$ (12)
To evaluate the accuracy of this ROM, we define a test parameter as $\boldsymbol{\mu}_{\_}{\text{test}}=[\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{1}{2},1]$, which lies within the parameter domain but was not included in the training set. The goal is to characterize the ROM error at this previously unseen parameter point, providing insight into the model’s generalization capabilities. Figure 2 shows the error between HDM and ROM displacement at the test parameter.
To characterize the error induced by ROM, we construct SROMs by replacing the deterministic ROB $\mathbf{V}$ with SROBs $\mathbf{W}$. The SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To achieve consistent UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the HDM displacement at training parameters, and $\mathbf{u}_{\_}L^{o}$ represents the ROM displacement at the same parameters. The optimal value of the hyperparameter $\beta$ is 20, which is optimized efficiently using the strategy given in Section 4. The hyperparameter training time for this example is 9.1 seconds.
Figure 3: Linear static problem: prediction by SS-PPCA.
Figure 3 shows the ROM error characterization for the linear parametric system at the test parameters. It can be observed that the SROM mean displacement is very close to the ROM displacement, and the SROM 95% PI is able to capture the ROM error consistently and sharply. The coverage for this example is 99%, that is, the SROM 95% PI has captured the ROM error over most of the domain. It can also be observed that this coverage is achieved while maintaining tight bounds around the HDM displacement.
6.2 Static problem: characterizing HDM error
In the previous example, we demonstrated the use of SROM to characterize the model truncation error induced by ROM in a linear parametric system. However, the utility of SROM is not limited to characterizing ROM error. It can also be employed to characterize discrepancies between HDM and experimental data, a crucial task in model validation. Here we characterize HDM error via the SROM approach in a linear static problem. The synthetic experimental data is generated by solving the system:
$\mathbf{K}_{\_}{\text{E}}\mathbf{x}_{\_}{\text{E}}=\mathbf{f},$ (13)
with $n$ (model dimension) =1,000absent1000=\$1,000\$= 1 , 000, $\mathbf{f}=\sum_{\_}{i=2}^{5}0.5\boldsymbol{\phi}^{i}+1\boldsymbol{\phi}^{6}$, and homogeneous Dirichlet boundary conditions $\mathbf{B}^{\intercal}\mathbf{x}_{\_}{\text{E}}=0$ where $\mathbf{B}=[\mathbf{e}_{\_}1\,\mathbf{e}_{\_}n]$. The stiffness matrix is defined as $\mathbf{K}_{\_}{\text{E}}=\mathbf{K}+\mathbf{K}_{\_}{\epsilon}$, where $\mathbf{K}$ is formulated the same as in the previous example and $\mathbf{K}_{\_}{\epsilon}$ is a perturbation matrix designed to induce a 5% change in the Euclidean norm of $\mathbf{K}_{\_}{\text{E}}$. In addition to the model error, we simulate measurement error by adding 5% noise to the experimental displacement $\mathbf{x}_{\_}{\text{E}}$. Hence, the observed experimental data follow: $\mathbf{x}_{\_}{\text{obs}}=\mathbf{x}_{\_}{\text{E}}+\eta$, where $\eta\sim N(0,\sigma^{2}\mathbf{I})$ with variance chosen to achieve the specified noise level. Experimental data can be very expensive to acquire. Therefore, we limit ourselves to experimental data from a sparse subset of spatial locations. These sparse measurements are used to characterize HDM error via the SROM approach. The computational HDM is defined as:
$\mathbf{K}\mathbf{x}_{\_}{\text{H}}=\mathbf{f},$ (14)
with the same $\mathbf{K}$ and $\mathbf{f}$ as above. However, the HDM produces biased predictions because the true system differs from this model.
To capture this error, we perturb the HDM so that the perturbed HDM responses can sufficiently cover the experimental data. The HDM is perturbed 100 times, and the responses are collected in a snapshot matrix. ROM of dimension $k=4$ is then constructed using POD, which further induces model truncation error. The ROM derived from the HDM is defined as:
$\mathbf{x}_{\_}{\text{R}}=\mathbf{V}\mathbf{q},\quad\mathbf{V}^{\intercal}\mathbf{K}\mathbf{V}\mathbf{q}=\mathbf{V}^{\intercal}\mathbf{f}.$ (15)
To characterize both the ROM and HDM error, we construct SROMs by substituting the deterministic ROB $\mathbf{V}$ with the SROBs $\mathbf{W}$. The SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. To have consistent UQ, the hyperparameter $\beta$ is optimized by minimizing the objective function given in Eq. 8. For this example, $\mathbf{u}_{\_}E$ represents the sparsely observable experimental displacement, and $\mathbf{u}_{\_}L^{o}$ denotes the ROM displacement at the same locations. The optimum value of the hyperparameter $\beta$ is 8, which is optimized efficiently using the strategy given in Section 4. The hyperparameter training time for this example is 1.9 seconds.
Figure 4: Linear static problem with experimental data: SS-PPCA prediction.
Figure 4 shows the HDM error characterization by the SS-PPCA method using the sparse experimental observations. It can be observed that our method successfully captures the model discrepancy, yielding accurate uncertainty quantification for the HDM predictions. The SROM 95% PI captures the noiseless ground-truth data well, and the bounds are very sharp. The trained SROM model via SS-PPCA enables fast prediction with error estimates and eliminates the dependence on computationally expensive HDM.
6.3 Dynamics problem: space structure
In Sections 6.1 and 6.2, we demonstrated that the SS-PPCA method performs well in characterizing uncertainty with low computational cost. However, those cases involved relatively simple linear static problems. To evaluate the robustness and scalability of the method, we compare it in a much more complex linear dynamics problem of a space structure. Because of the higher model complexity and more realistic structural dynamics, this example may better assess the efficiency and effectiveness of the method.
Figure 5: Dynamics problem: (a) space structure; (b) loading.
We consider a major component of a space structure [45] given an impulse load in the $z$-direction, see Figure 5. The space structure consists of two major parts: an open upper and a lower part. A large mass (approximately 100 times heavier than the remainder of the model) is placed at the center of the upper part. This mass is connected to the sidewalls via rigid links. The lower part of the structure consists of the outer cylindrical shell and an inner shock absorption block with a hollow space to hold the essential components. The impulse load, as shown in Figure 5b, is applied at the center of the mass in the $z$-direction, which is transmitted to the upper part through the sidewalls and then to the lower part via the mounting pedestal. The impulse load can compromise the functionality of the essential components located in the lower part. Hence, monitoring the acceleration and velocity at the critical points of the essential parts is important for their safe operation. We take the quantity of interest (QoI) to be the $x$-velocity of a critical point at one of the essential components. The space structure has no boundary conditions; it can be assumed to be floating in outer space. The governing equation of the HDM of the space structure is given by:
$\mathbf{M}_{\_}{\text{H}}\,\ddot{\mathbf{x}}+\mathbf{C}_{\_}{\text{H}}\,\dot{\mathbf{x}}+\mathbf{K}_{\_}{\text{H}}\,\mathbf{x}=\mathbf{f},$ (16)
[…]
$\mathbf{x}_{\_}{\text{R}}=\mathbf{V}\mathbf{q},\quad\mathbf{M}_{\_}{\text{H},\mathbf{V}}\,\ddot{\mathbf{q}}+\mathbf{C}_{\_}{\text{H},\mathbf{V}}\,\dot{\mathbf{q}}+\mathbf{K}_{\_}{\text{H},\mathbf{V}}\,\mathbf{q}=\mathbf{V}^{\intercal}\mathbf{f},$ (17)
where reduced matrices are denoted with a pattern $\mathbf{A}_{\_}{\mathbf{V}}:=\mathbf{V}^{\intercal}\mathbf{A}\mathbf{V}$, so that $\mathbf{M}_{\_}{\text{H},\mathbf{V}}=\mathbf{V}^{\intercal}\mathbf{M}_{\_}{\text{H}}\mathbf{V}$, for example. The initial condition of ROM is given by $\mathbf{q}(0)=0$ and $\dot{\mathbf{q}}(0)=0$. In contrast to the high simulation time of HDM, ROM takes 0.2 seconds—roughly 11,200 times faster. However, this computational efficiency comes at the expense of reduced accuracy. Hence, error characterization of ROM is essential, which is done by SROM throughout this study. To construct the SROMs, we substitute the deterministic ROB $\mathbf{V}$ by the SROBs $\mathbf{W}$.
For the SS-PPCA method, the SROBs $\mathbf{W}$ are sampled using the stochastic subspace model described in Section 3.3 and Algorithm 1. For this example, $\mathbf{u}_{\_}E$ corresponds to the velocity from the HDM at a critical structural node, and $\mathbf{u}_{\_}{L}^{o}$ corresponds to the velocity predicted by ROM at the same location. The optimum value of the hyperparameter $\beta$ comes out to be 39, which is optimized using the strategy given in Section 4. The hyperparameter training time for this example is 785 seconds, which is low compared to the computational cost of the HDM. This shows that our method is scalable and can provide fast and reliable predictions even in complex systems.
Figure 6: Dynamic prediction of the SS-PPCA model: velocity.
The HDM velocity shows a high-frequency behavior, especially during the initial 10 milliseconds. The ROM with just 10 modes fails to capture this behavior. Figure 6 illustrates that the 95% PI of the SS-PPCA method is notably consistent and significantly sharp, even in the initial 10 milliseconds period (see the Figure 6 inset for $[0,10]$ ms). The SS-PPCA method not only captures the high-frequency behavior but also maintains tight bounds around it. Additionally, the SROM mean velocity of SS-PPCA closely matches the deterministic ROM velocity, as it should be. The coverage for velocity via the SS-PPCA method is 94.83%.
Figure 7: Dynamic prediction of the SS-PPCA model: acceleration.
Figure 8: Dynamic prediction of the SS-PPCA model: displacement.
Figure 7 shows the $x$-direction acceleration at the critical node predicted by the SS-PPCA. The acceleration also exhibits a high-frequency behavior, due to the impulse loading. Despite this, our SS-PPCA method is able to produce consistent UQ with sharp bounds. It is important to note that even though velocity (the QoI for this example) is used for training the hyperparameters and displacement is used for obtaining the deterministic ROB $\mathbf{V}$, our approach enables predictions of unobserved quantities such as acceleration. The fact that acceleration has not been observed makes our result very intriguing. Furthermore, our approach enables efficient predictions of any unobserved QoIs of the whole system, a capability that many studies in the literature fail to achieve. This highlights the effectiveness and practicality of our approach to engineering applications. The coverage for acceleration via the SS-PPCA method is 85.49%, which means that the error characterization can be improved a bit. However, given that the acceleration is an unobserved quantity and exhibits a high-frequency behavior, the result is quite good.
Figure 8 compares the $x$-direction displacement at the critical node predicted by the SS-PPCA. It can be observed that SS-PPCA displacement prediction is highly accurate: the SROM, ROM, and HDM show similar behavior, and the 95% PI is very tight. The coverage for displacement via the SS-PPCA method is 96.78 %, which means that the displacement prediction is consistent and highly accurate.
Figure 9: Dynamic prediction of the SS-PPCA model: velocity at a random DoF.
All the quantities compared in Figures 6, 7 and 8, whether observed or unobserved, are from the same node of the space structure. However, we may need to make a prediction at a different node of the system, or we may be interested in the general behavior of the system. Figure 9 shows the prediction of the SS-PPCA method for the $y$-direction velocity at a random node; the result is still consistent and very sharp. This shows that the SS-PPCA methods allow us to make predictions about the entire structure just by training on the observation data of a single DoF.
In summary, the low computational cost of hyperparameter optimization of SS-PPCA, combined with the low computational cost of ROM, can accelerate the prediction of complex engineering phenomena. Furthermore, SS-PPCA offers a fully automated, efficient, and principled training workflow with minimal user input. This makes it well suited for practical applications. One issue with methods based on SROM is that they cannot fully eliminate model error as they rely on a ROM. This issue can be addressed by increasing the number of reduced bases or incorporating model closure terms. A study on model closure for ROMs is presented by Ahmed et al. [46].
7 Conclusion
We introduced a novel stochastic subspace model, which is used to characterize model error in the framework of stochastic reduced-order models. It is simple, efficient, easy to implement, and well-supported by analytical results. Our SS-PPCA model has only one hyperparameter, which is optimized systematically to improve consistency. Through various numerical examples, we reveal the characteristics of this model, show its flexibility across single and multiple model cases, establish its consistency, and quantify its remarkable sharpness. It opens up a promising path to the challenging problem of model-form uncertainty.
Across all examples, the stochastic subspace model effectively captures various forms of model error, including parametric uncertainty, ROM error, and HDM errors informed by noisy experimental data. Although the examples used in this study are linear systems, our method is generally applicable for characterizing model error in nonlinear systems. However, in scenarios with large errors, mere characterization may not suffice. In such cases, future work will focus on developing practical and efficient techniques for model error correction.
Acknowledgments
The authors thank Johann Guilleminot and Christian Soize for invaluable discussions on this topic. This work was supported by the University of Houston.
Appendix A Mode of $\text{MACG}_{\_}{n,k}(\boldsymbol{\Sigma})$
In this section, we establish the mode of the MACG probability distribution on Grassmann manifolds. To the best of our knowledge, this is the first appearance of such results.
The matrix angular central Gaussian distribution $\text{MACG}_{\_}{n,k}(\boldsymbol{\Sigma})$ on the Grassmann manifold $\text{Gr}(n,k)$ [44, 47] is the probability distribution with the probability density function (PDF) $p(\mathbf{X}\mathbf{X}^{\intercal};\boldsymbol{\Sigma})=z^{-1}|\mathbf{X}^{\intercal}\boldsymbol{\Sigma}^{-1}\mathbf{X}|^{-n/2}$ with respect to the normalized invariant measure of $\text{Gr}(n,k)$. Here $\boldsymbol{\Sigma}>0$ is an order-$n$ symmetric positive definite matrix, a $k$-dim subspace $\mathscr{X}=\text{range}(\mathbf{X})$ is represented by the orthogonal projection matrix $\mathbf{X}\mathbf{X}^{\intercal}$ where $\mathbf{X}\in\text{St}(n,k)$ is an orthonormal basis, and the normalizing constant $z=|\boldsymbol{\Sigma}|^{k/2}$ where ||$\cdot$|| ⋅ | denotes the matrix determinant. In the following, let $\boldsymbol{\Sigma}$ have eigenvalues $\lambda_{\_}1\geq\cdots\geq\lambda_{\_}n>0$ and corresponding eigenvectors $\mathbf{v}_{\_}1,\cdots,\mathbf{v}_{\_}n$ that form an orthonormal basis of $\mathbb{R}^{n}$.
Theorem 1.
The PDF of the $\text{MACG}_{\_}{n,k}(\boldsymbol{\Sigma})$ distribution on $\text{Gr}(n,k)$ is maximized at any $k$-dim principal subspace of $\boldsymbol{\Sigma}$. If $\lambda_{\_}k>\lambda_{\_}{k+1}$, then the principal subspace $\mathscr{V}_{\_}k=\text{range}(\mathbf{V}_{\_}k)$ with $\mathbf{V}_{\_}k=[\mathbf{v}_{\_}1\cdots\mathbf{v}_{\_}k]$ is the unique mode of the distribution.
Proof.
Let us first consider a special case when $\boldsymbol{\Sigma}$ is the diagonal matrix $\boldsymbol{\Lambda}=\operatorname{diag}(\lambda_{\_}1,\cdots,\lambda_{\_}n)$. Let subspace $\mathscr{X}\sim\text{MACG}_{\_}{n,k}(\boldsymbol{\Lambda})$ and have an orthonormal basis $\mathbf{X}\in\text{St}(n,k)$. The PDF can be written as:
$p(\mathscr{X})=|\boldsymbol{\Lambda}|^{-k/2}|\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}|^{-n/2}=\left(\prod_{\_}{i=1}^{n}\lambda_{\_}i\right)^{-k/2}|\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}|^{-n/2}.$ (18)
We see that the PDF is a smooth function of $\mathbf{X}$, and we first characterize its critical points. Taking the logarithm of the above equation, we get:
$\log p(\mathscr{X})=-\frac{n}{2}\log|\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}|-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i.$ (19)
The second term $\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i$ is a constant, so the critical points of $p(\mathscr{X})$ are the critical points of $\log|\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}|$, which can be computed by setting $\operatorname{\text{tr}}(\mathbf{S}^{-1}\partial\mathbf{S})=0$ where $\mathbf{S}=\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}$. The partial derivative of $\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}$ can be written as $\partial(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})=(\partial\mathbf{X})^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}+\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}(\partial\mathbf{X})$. The differential $\partial\mathbf{X}\in T_{\_}{\mathbf{X}}\text{St}(n,k)$, meaning that it lies in the tangent space of the Stifel manifold $\text{St}(n,k)$ at the point $\mathbf{X}$, which must satisfy $(\partial\mathbf{X})^{\intercal}\mathbf{X}+\mathbf{X}^{\intercal}(\partial\mathbf{X})=0$. In other words, $(\partial\mathbf{X})^{\intercal}\mathbf{X}$ must be antisymmetric, which allows us to write $\partial\mathbf{X}$ in a projective form: $\partial\mathbf{X}=\mathbf{M}-\mathbf{X}\operatorname{\text{sym}}(\mathbf{X}^{\intercal}\mathbf{M})$ where $\mathbf{M}\in\mathbb{R}^{n\times k}$ and $\text{sym}(\mathbf{A})=(\mathbf{A}+\mathbf{A}^{\intercal})/2$. Now we have:
$\operatorname{\text{tr}}(\mathbf{S}^{-1}\partial\mathbf{S})$ $=\operatorname{\text{tr}}\{(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}[(\partial\mathbf{X})^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}+\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}(\partial\mathbf{X})]\}$ (20)
$=\operatorname{\text{tr}}\{(\partial\mathbf{X})^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X}(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}+(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}(\partial\mathbf{X})\}$
$=2\operatorname{\text{tr}}\{(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}(\partial\mathbf{X})\}$
$=2\operatorname{\text{tr}}\{(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}[\mathbf{M}-\mathbf{X}\operatorname{\text{sym}}(\mathbf{X}^{\intercal}\mathbf{M})]\}$
$=2\operatorname{\text{tr}}\{(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{M}-\operatorname{\text{sym}}(\mathbf{X}^{\intercal}\mathbf{M})\}$
$=2\operatorname{\text{tr}}\{(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{M}-\mathbf{X}^{\intercal}\mathbf{M}\}$
$=2\,\text{sum}\{[(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}-\mathbf{X}^{\intercal}]\circ\mathbf{M}\}.$
With $\operatorname{\text{tr}}(\mathbf{S}^{-1}\partial\mathbf{S})=0$ for all $\mathbf{M}\in\mathbb{R}^{n\times k}$, we have $(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}-\mathbf{X}^{\intercal}=0$, that is, $\mathbf{X}=\boldsymbol{\Lambda}^{-1}\mathbf{X}(\mathbf{X}^{\intercal}\boldsymbol{\Lambda}^{-1}\mathbf{X})^{-1}$. Further simplification gives us $(\mathbf{X}\mathbf{X}^{\intercal})(\boldsymbol{\Lambda}^{-1}\mathbf{X})=(\boldsymbol{\Lambda}^{-1}\mathbf{X})$. Because $\mathbf{X}\mathbf{X}^{\intercal}$ is the orthogonal projection onto $\text{range}(\mathbf{X})$, we have $\text{range}(\boldsymbol{\Lambda}^{-1}\mathbf{X})\subseteq\text{range}(\mathbf{X})$; since both sides of the equation are $k$-dim subspaces, we have $\text{range}(\boldsymbol{\Lambda}^{-1}\mathbf{X})=\text{range}(\mathbf{X})$. Therefore, $\mathscr{X}=\text{range}(\mathbf{X})$ is an invariant $k$-subspace of $\boldsymbol{\Lambda}^{-1}$, or equivalently, an invariant $k$-subspace of $\boldsymbol{\Lambda}$. If the eigenvalues $(\lambda_{\_}i)_{\_}{i=1}^{n}$ are distinct, then $\mathscr{X}$ must contain a $k$-combination of the standard basis $\{\mathbf{e}_{\_}i\}_{\_}{i=1}^{n}$.
So far we have proved that the critical points of $p(\mathscr{X})$ are the invariant $k$-subspaces of $\boldsymbol{\Lambda}$. Since the Grassmann manifold $\text{Gr}(n,k)$ is compact and without a boundary, the smooth PDF $p(\mathscr{X})$ has a global minimum and a global maximum, both of which are critical points. We define two $k$-dim subspaces ¯_$\overline{\mathscr{E}}_{\_}k$ and ¯_$\underline{\mathscr{E}}_{\_}k$ as the ranges of the bases $\overline{\mathbf{E}}_{\_}k:=[\mathbf{e}_{\_}1\cdots\mathbf{e}_{\_}k]$ and $\underline{\mathbf{E}}_{\_}k:=[\mathbf{e}_{\_}{n-k+1}\cdots\mathbf{e}_{\_}n]$, respectively. We have:
$\log p(\overline{\mathscr{E}}_{\_}k)$ $=-\frac{n}{2}\log|\overline{\mathbf{E}}_{\_}{k}^{\intercal}\boldsymbol{\Lambda}^{-1}\overline{\mathbf{E}}_{\_}k|-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i$ (21)
$=-\frac{n}{2}\log|\operatorname{diag}(\lambda_{\_}1^{-1},\cdots,\lambda_{\_}k^{-1})|-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i$
$=-\frac{n}{2}\sum_{\_}{i=1}^{k}\log\lambda_{\_}{i}^{-1}-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i$
$=\frac{n}{2}\sum_{\_}{i=1}^{k}\log\lambda_{\_}{i}-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i.$
Similarly, we have $\log p(\underline{\mathscr{E}}_{\_}k)=\frac{n}{2}\sum_{\_}{i=n-k+1}^{n}\log\lambda_{\_}{i}-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i$. In general, let $\lambda_{\_}{a_{\_}1}\geq\cdots\geq\lambda_{\_}{a_{\_}k}$ be the eigenvalues associated with an invariant $k$-subspace $\mathscr{X}_{\_}0$, then:
$\log p(\mathscr{X}_{\_}0)=\frac{n}{2}\sum_{\_}{j=1}^{k}\log\lambda_{\_}{a_{\_}j}-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i.$ (22)
Since $\lambda_{\_}1\geq\cdots\geq\lambda_{\_}n$, we have $p(\underline{\mathscr{E}}_{\_}k)\leq p(\mathscr{X}_{\_}0)\leq p(\overline{\mathscr{E}}_{\_}k)$, for all critical points $\mathscr{X}_{\_}0$. This means that ¯_$\underline{\mathscr{E}}_{\_}k$ and ¯_$\overline{\mathscr{E}}_{\_}k$ are the global minimal point and the global maximal point of $p(\mathscr{X})$, respectively; that is, $p(\underline{\mathscr{E}}_{\_}k)\leq p(\mathscr{X})\leq p(\overline{\mathscr{E}}_{\_}k)$.
Now we show that a critical point $\mathscr{X}_{\_}0$ cannot be a local maximum unless $\{\lambda_{\_}{a_{\_}1},\cdots,\lambda_{\_}{a_{\_}k}\}=\{\lambda_{\_}1,\cdots,\lambda_{\_}k\}$. Let $\mathbf{v}_{\_}{a_{\_}1},\cdots,\mathbf{v}_{\_}{a_{\_}k}$ be eigenvectors associated with eigenvalues $\lambda_{\_}{a_{\_}1},\cdots,\lambda_{\_}{a_{\_}k}$, respectively, such that they form an orthonormal basis of $\mathscr{X}_{\_}0$. Assume that eigenvalue $\lambda^{\prime}\notin\{\lambda_{\_}{a_{\_}1},\cdots,\lambda_{\_}{a_{\_}k}\}$ and $\lambda^{\prime}>\lambda_{\_}{a_{\_}k}$. Let $\mathbf{v}^{\prime}$ be an eigenvector associated with the eigenvalue $\lambda^{\prime}$. Consider the trajectory $\mathscr{X}(\theta)=\text{range}(\mathbf{X}(\theta))$, where $\mathbf{X}(\theta)=[\mathbf{v}_{\_}{a_{\_}1}\,\cdots\,\mathbf{v}_{\_}{a_{\_}{k-1}}\,\mathbf{v}(\theta)]$, $\mathbf{v}(\theta)=\cos(\theta)\mathbf{v}_{\_}{a_{\_}k}+\sin(\theta)\mathbf{v}^{\prime}$, and $\theta\in[0,\frac{\pi}{2}]$. It is a constant-speed rotation that starts from $\mathscr{X}_{\_}0$. We can show that:
$\log p(\mathscr{X}(\theta))=\frac{n}{2}\left(\sum_{\_}{j=1}^{k-1}\log\lambda_{\_}{a_{\_}j}-\log\left(\lambda_{\_}{a_{\_}k}^{-1}\cos^{2}\theta+\lambda^{\prime-1}\sin^{2}\theta\right)\right)-\frac{k}{2}\sum_{\_}{i=1}^{n}\log\lambda_{\_}i,$ (23)
which increases monotonically. Therefore, $\mathscr{X}_{\_}0$ is not a local maximum.
In summary, $\text{MACG}_{\_}{n,k}(\boldsymbol{\Lambda})$ has its global maximum at ¯_$\overline{\mathscr{E}}_{\_}k$, or any $k$-dim principal subspace of $\boldsymbol{\Lambda}$. These subspaces are also the modes of the distribution. If $\lambda_{\_}k>\lambda_{\_}{k+1}$, there is a unique $k$-dim principal subspace ¯_$\overline{\mathscr{E}}_{\_}k$, which is the unique mode of the distribution.
For a general $\boldsymbol{\Sigma}$, it has an eigenvalue decomposition $\boldsymbol{\Sigma}=\mathbf{V}\boldsymbol{\Lambda}\mathbf{V}^{\intercal}$ where $\mathbf{V}=[\mathbf{v}_{\_}1\,\cdots\mathbf{v}_{\_}k]$. Given a change of basis $\mathbf{x}=\mathbf{V}\mathbf{z}$, the previous arguments still hold, and the theorem follows immediately. ∎